OpenStudy (anonymous):

Evaluate 4x-3,ifx=2 1/4

OpenStudy (abb0t):

2\(\sf \frac{1}{4}\) = 2.25

OpenStudy (abb0t):

That means \(\sf \color{red}{x=2.25}\), then 4\(\sf \color{red}{x}\)-3

OpenStudy (anonymous):

6

OpenStudy (abb0t):

maddie, giving the answer doesn't help the person. Help is what this site is for.

OpenStudy (anonymous):

you need to fill all the xs in for the problem, so you have...4(2.25)-3 4*2.25=9 9-3=6